Wath is the anser to x/3+1/2=4
Answer by checkley71(8403)

X/3+1/2=4 OR ADDING THE TWO FRACTION WITH A COMMON DENOMINATOR OF 6 WE GET
(2X+3)/6=4 CROSS MULTIPLYING WE GET 2X+3=24 OR 2X=21 OR X=10.5
PROOF 10.5/3+1/2=4 OR 3.5+.5=4.